OpenStudy (anonymous):

if 5 out of 25 people like chocolate ice cream what percent of the poeple do NOT like chocolate ice cream choices first to give me the answer i will fan give best response 20% 5% 80% 30% plz just tell me the answer

OpenStudy (anonymous):

5-25=20 to get the percentage 20/25=80%

OpenStudy (anonymous):

@This_Is_Batman @bibby @_mar5iah_ @

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

5/25=.2 So 20% of the people like it. Total is 100%, so 100-20=80% :DDD

OpenStudy (anonymous):

thx bats

OpenStudy (anonymous):

No problem.:p